# Gravity from thermodynamics in vacuum: Lorentz invariance and the role of Bel-Robinson tensor

### v1

@@ Abstract
Thermodynamics of local causal horizons have been shown to encode the information necessary to derive the equations governing the gravitational dynamics. We have previously shown that, in the presence of matter, this derivation further implies quantum phenomenological corrections to gravitational dynamics. Herein, we study whether similar corrections also occur in vacuum. We show that, under the assumptions of locality and local Lorentz invariance of physics, the vacuum dynamics is prescribed by the Einstein equations. We also discuss an alternative paradigm which assumes the existence of a preferred direction of time (much like in Einstein-aether or Hořava-Lifshitz gravity). Then, we find an modified gravitational dynamics in which Ricci curvature is sourced by Bel-Robinson super-energy.
@@ I Introduction
The black hole entropy not only follows directly from the gravitational Lagrangian, but actually contains enough information to reconstruct the gravitational dynamics [4, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. More precisely, the dynamics can be derived by first requiring that gravitational entropy is not specific for black holes, but instead a universal property of any causal horizon. This assumption is supported by a number of results obtained for thermodynamics of various horizons, see e.g. [26, 27, 28, 29, 30, 31, 32, 25]. Then, one can construct a local, observer-dependent causal horizon in every regular spacetime point and study the changes in its entropy. It turns out that the thermodynamic equilibrium conditions imposed on such local horizons encode (locally) the equations governing the gravitational dynamics. If the equivalence principle [^1] More precisely, one in general needs the Einstein equivalence principle. Specifically, to recover the Einstein equations a more stringent formulation, the strong equivalence principle (valid for gravitational test physics) is required [15]. For detailed discussion of the various formulations of the equivalence principle and their relations, see [33]. holds, these equations are valid throughout the spacetime. This approach not only recovers Einstein equations [4, 15, 19], but also the equations of motion of any gravitational theory whose Lagrangian is an arbitrary function of the metric and the Riemann tensor [16, 20, 21, 22], using the corresponding gravitational Wald entropy expression [6]. Moreover, these local equilibrium conditions phrased instead in terms of quantum von Neumann entropy have been shown to encode the semiclassical Einstein equations in which the quantum expectation value of the energy-momentum tensor serves as the source term for the classical spacetime curvature [17, 18, 19, 20].
@@
Herein, inspired by this success of thermodynamic methods in recovering the classical and semiclassical gravitational dynamics, we, following a research program we started some years ago, employ them to peer one step further, to low energy quantum gravitational effects. To this end, we consider the leading order quantum gravitational correction to entropy of causal horizon, a term proportional to the logarithm of the horizon area. Remarkably, a leading order correction term of this form is nearly universally predicted by various approaches to quantum gravity [34, 35, 36, 37, 38, 39] and also by some theory-independent considerations [40, 41, 42, 43, 44, 45, 46, 47, 48]. This generality of the logarithmic term in horizon entropy makes its influence on gravity robust and independent of any specific approach to quantum gravity. We have previously studied the effect of this term on the linearised gravitational dynamics, finding a result equivalent to the linearised equations of motion of quadratic gravity. We have also shown that beyond the linearised regime, it leads to correction terms quadratic in the Ricci tensor which do not appear in quadratic gravity, marking a clear departure from it. However, in both special cases we studied so far the corrections to gravitational dynamics that only occur in the presence of matter. In the present work, we ask whether the Einstein equations are modified even in vacuum. We show that imposing the local Lorentz invariance prohibits any such modifications. Nevertheless, if one chooses to introduce a preferred local direction of time, the thermodynamic derivation yields a scalar equation resembling the Hamiltonian constraint of general relativity with a source term proportional to the Bel-Robinson tensor contracted with the local direction of time in all indices [49, 50] (such a scalar is known as the Bel-Robinson super-energy density). The Bel-Robinson tensor is quadratic in the Weyl tensor and satisfies an analogy of the dominant energy condition valid for the energy-momentum tensor. It has been suggested that the Bel-Robinson super-energy density can be interpreted as a quasilocal gravitational energy density per unit surface area [51, 50]. This interpretation appears to be consistent with the role the Bel-Robinson super-energy density plays in the geometry of small spheres [52, 53]. Therefore, its appearance as a source term for the Hamiltonian constraint is somewhat intuitive, since in general relativity the source term is the matter-energy density.
@@ II.1 Light-cone cut local causal diamonds
We start by introducing the specific realisation of local, observer-dependent causal horizons we work with in this paper: the causal diamonds. In flat spacetime a causal diamond can be simply defined as the domain of dependence of a spacelike co-dimension $1$ ball. In a generic curved spacetime only small causal diamonds can be meaningfully defined. More specifically, one requires that the geodesic radius of the co-dimension $1$ ball forming the base of the causal diamond is much smaller than the local curvature length scale (i.e. the inverse of the square root of the largest eigenvalue of the Riemann tensor). Then, the spacetime curvature effects on the geometry of the causal diamonds can be treated as small corrections. However, even in this case, there exist several inequivalent generalisations of a causal diamond [55]. While we could in principle consider any of them, the one best suited for our purposes is the light-cone cut local causal diamond due to the way it is defined, as we explain in the following.
The construction of a light-cone cut local causal diamond starts at an arbitrary regular spacetime point $A_{\text{p}}$, the past apex of the eventual causal diamond. At this point, we choose a unit timelike vector $n^{\mu}$ to represent a local direction of time. We then consider a class of future directed null vectors $k^{\mu}$ at $A_{\text{p}}$ normalised so that $n_{\mu}k^{\mu}=-1$, and define the past boundary of the light-cone cut local causal diamond as the congruence of null wordlines tangent to $k^{\mu}$. Any spacelike cross-section of this congruence is an approximate $2$-sphere (up to curvature-dependent corrections) and its interior an approximate $3$-dimensional spacelike ball. The base of the causal diamond corresponds to its spatial slice $\Sigma_{0}$ at the affine geodesic parameter length $l$. To construct the future (contracting) part of the light-cone cut local causal diamond, we introduce a future-directed null vector field $l^{\mu}$ defined on the boundary of $\Sigma_{0}$, so that $n_{\mu}l^{\mu}=-1$. Furthermore, denoting the projection of $k^{\mu}$ on the surface orthogonal to $n^{\mu}$ by $m^{\mu}$ (this vector is simply a unit, outward-pointing radial vector), we choose $l^{\mu}$ to have this projection equal to $-m^{\mu}$. Then, the contracting congruence of the null geodesics tangent to $l^{\mu}$ makes up the future boundary of the causal diamond. We show a sketch of the light-cone cut local causal diamond in figure 1.
@@
The light-cone cut local causal diamonds then offer the additional advantage of fully specifying the null boundary of the causal diamond (up to some residual freedom [22, 25], we discuss in the following), while keeping the geometry of its spatial cross-sections somewhat ambiguous [55]. Then, it becomes very simple to track the physical matter entropy flux across the boundary and the corresponding changes in its area. Since, as we review in the following, the equilibrium conditions imposed on these two processes encode the equations governing the gravitational dynamics, light-cone cut local causal diamonds are ideally suited for deriving them in this way.
@@ II.2 Gravitational dynamics from thermodynamics
We have seen that the null boundary of a light-cone cut causal diamond represents a conformal Killing horizon. Therefore, it can be argued that it possesses entropy proportional to the area $\mathcal{A}$ of its spatial cross-section, i.e., $S=\eta\mathcal{A}$, where $\eta$ denotes the proportionality constant (see [25] and references therein for a detailed justification of this assumption). For the purposes of the present work, we do not care about the microscopic origin of this entropy, we simply take it as our assumption. The entropy, of course, changes with the expansion of the horizon. At the same time, matter flux crossing the horizon also carries entropy with it. It has been shown that if the expansion of the causal diamond corresponds to a reversible thermodynamic process, the balance between both entropies encodes the gravitational dynamics [22, 25]. In the following, we review the main points of this argument, first recalling the expressions for the changes of both relevant entropies.
To compute the matter entropy flux, we employ the Clausius equilibrium relation $\Delta S_{\text{C}}=\Delta Q/T_{\text{U}}$, where $\Delta Q$ denotes the heat flux across the horizon and $T_{\text{U}}$ the Unruh temperature measured by accelerating observers moving inside the causal diamond. The total flux of this Clausius entropy over the past boundary of the causal diamond equals (for details on its definition and computation, see [57])
@@
For simplicity, we set the expansion and the shear to vanish at the base of the light-cone cut causal diamond [^2] In principle, one should also be able to carry out the derivation with a generic $\sigma_{0,\mu\nu}$, which should correspond to non-equilibrium entropy production [15]. We expect to address this generalisation in a future work., i.e., $\theta_{0}=\sigma_{0\mu\nu}=0$. Then, it is easy to see that the shear contributions to the expansion are not relevant and we obtain the following solution
@@
Integrating, we finally obtain
@@
Equations (16) closely resemble the traceless Einstein equations. By taking their Newtonian limit, we can further define the Newton gravitational constant $G$ in terms of the Planck constant $\hbar$ and the entropy proportionality constant $\eta$, i.e., $G=1/\left(4\hbar\eta\right)$. Then, we indeed recover the traceless Einstein equations
@@
Let us conclude with two short remarks regarding the interpretation of equations (17). First, they are traceless, with the cosmological constant appearing only as an on-shell integration constant. Moreover, the geometry of the light-cone cut local causal diamonds is insensitive to the overall conformal factor of the metric. We have previously shown that these features imply that the gravitational dynamics we recover from thermodynamics (without any additional assumptions) does not correspond to general relativity, but rather to Weyl transverse (unimodular) gravity [23, 24, 25]. Therefore, the corrections to gravitational dynamics we discuss in the following should also be seen as modifying Weyl transverse gravity rather than general relativity.
Second, the fact that thermodynamics of local causal horizons, together with the equivalence principle, suffices to recover the equations governing gravitational dynamics is often interpreted as suggesting that gravity is emergent. In other words, the Einstein equations would be essentially equations of state describing the behaviour of some unknown quantum degrees of freedom of the spacetime in the classical limit [4, 14, 15, 59]. While we find this viewpoint worthwhile and thought provoking, we believe it is simply unnecessary to embrace it. We have shown that thermodynamics leads to novel insights regarding gravitational dynamics without any relation to its emergence [23, 24, 25, 60]. All we need to assume is that the local equilibrium conditions encode all the information necessary to reconstruct the gravitational dynamics. We maintain this viewpoint in the rest of the paper.
@@ III Higher order corrections in vacuum: the Bel-Robinson tensor
In the previous section, we have reviewed the derivation of the traceless Einstein equations from thermodynamics. The derivation neglected the subleading contributions to the curved spacetime changes of the area of the spatial cross-section of the horizon, and, hence, to its entropy. It is natural to ask whether these terms encode some corrections to Einstein equations. In some special cases, we have previously shown that this is indeed the case, although such corrections only occur if we also take into account the subleading quantum contribution to the horizon entropy proportional to the logarithm of its area [60, 61]. Specifically, we have shown that the derivation generically leads to corrections to the traceless Einstein equations quadratic in the Ricci tensor [61, 62]. We have also derived the linearised modified dynamics, finding a result equivalent to a special case of linearised quadratic gravity [60]. Due to various technical and conceptual challenges, we are for the time being unable to derive the modified equations in the fully general case. Here, we instead study another simplified setting, the vacuum case. On the one hand, we use this setting to develop tools that should allow us to derive the completely general equations in a subsequent work. On the other hand, any possible modifications to vacuum gravitational dynamics are of great interest by themselves, being the only kind of corrections relevant to the standard black hole solutions, in particular, to the Kerr metric (see [63, 64] for examples of such corrections). They are also important for the behaviour of the spacetime in the vicinity of generic spacelike singularities, as described by the Belinskii-Khalatnikov-Lifshitz conjecture [65]. Finally, such corrections generically modify the propagation of gravitational waves, which might make them experimentally falsifiable by the current or future gravitational waves detectors [66]. To study possible vacuum corrections to gravitational dynamics implied by the thermodynamic derivation, we can follow the same basic strategy we employed to derive the traceless Einstein equations. However, we need to introduce some conceptual changes, both on the level of the initial assumptions and in the process of the derivation. Therefore, we start by reviewing the requirements we impose.
@@
In the semiclassical setup, we have relied on the Einstein equivalence principle to construct the local Minkowski vacuum necessary for invoking the Unruh effect. However, the status of any formulation of the equivalence principle in a regime in which quantum gravitational corrections become relevant remains an open issue. Therefore, we now instead impose two more modest requirements, which do not rely on the general formulation of the equivalence principle but still allow for a thermodynamics description of local causal horizons. Our first requirement is that a detector which moves with an approximately constant acceleration $a$ for a sufficient time to thermalise, sees the local Minkowski vacuum as a thermal bath of particles at the Unruh temperature $T_{\text{U}}=a/\left(2\pi\right)$, provided that the spacetime curvature effects on the detector can be neglected. While the fate of the Unruh effect in quantum gravity remains an open questions, there exist indications that it remains viable [67].
@@
The second requirement we impose is that entropy of a spherically symmetric local causal horizon in four spacetime dimensions has a generic form
@@
where $\mathcal{A}_{0}$ is a constant with the dimensions of area and $\mathcal{C}$ a real number. The first two terms remain in the limit $l_{\text{P}}\to 0$ (and the leading one diverges), whereas the $O\left(\mathcal{A}_{0}/\mathcal{A}\right)$ contributions vanish. Then, restricting our attention to the area term and the logarithm is equivalent to keeping the only two contributions to entropy relevant for large horizons, such that $\mathcal{A}\gg l_{\text{P}}^{2}$.
Equation (18) arises independently in a number of different contexts. It appears, e.g. in loop quantum gravity [34, 35], string theory [36, 37, 38], AdS/CFT correspondence [39], entanglement entropy calculations [30, 41], phenomenological approaches based on the existence of a minimal length [42, 43], in an analysis of statistical fluctuations [44, 45, 46] and from quantisation of the horizon area [47, 48]. In other words, the logarithmic correction to entropy associated with causal horizons is present in most of the approaches to quantum gravity, where the specific approach is characterised just by the sign and value of the parameter $\mathcal{C}$. Thence, the modifications of gravitational dynamics implied by the logarithmic term are in principle relevant for any approach to quantum gravity and are only parametrised by a single unknown number, $\mathcal{C}$, which is of the order of unity for most approaches.
@@
We now need to solve the Raychaudhuri equation (5) and equation (10) for the evolution of the shear tensor to the order $O\left(\lambda^{3}\right)$ (recall that the affine null geodesic parameter $\lambda$ lies in the interval from $-l$ to 00 on the past boundary), which becomes relevant once we take into account the logarithmic corrections to entropy. We obtain (keeping in place the condition $\theta_{0}=\sigma_{0\mu\nu}=0$)
$\theta\left(x\right)=$ $\theta_{\text{flat}}\bigg{[}1+\frac{\lambda^{2}}{1+\lambda\theta_{\text{flat}}}R_{\mu\nu}k^{\mu}k^{\nu}+\frac{1}{6}\frac{\lambda^{5}\theta_{\text{flat}}}{\left(1+\lambda\theta_{\text{flat}}\right)^{2}}\lambda^{3}\left(R_{\mu\nu}k^{\mu}k^{\nu}\right)^{2}\bigg{]}$

### v2

@@ Abstract
Thermodynamics of local causal horizons has been shown to encode the information necessary to derive the equations governing the gravitational dynamics. We have previously argued that, in the presence of matter, this derivation further implies quantum phenomenological corrections to gravitational dynamics. Herein, we study whether similar corrections also occur in vacuum. We show that, under the assumptions of locality and local Lorentz invariance of physics, the vacuum dynamics is prescribed by the Einstein equations. We also discuss an alternative paradigm which assumes the existence of a preferred direction of time (much like in Einstein-aether or Hořava-Lifshitz gravity). Then, we find an modified gravitational dynamics in which Ricci curvature is sourced by Bel-Robinson super-energy.
@@ I Introduction
The black hole entropy not only follows directly from the gravitational Lagrangian, but actually contains enough information to reconstruct the gravitational dynamics [4, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. It requires extending the gravitational entropy from being specific to black holes to a universal property of any causal horizon. This requirement is supported by the analysis of thermodynamics of various horizons, including de Sitter, cosmological and acceleration horizons [26, 27, 28, 29, 30, 31, 32, 25]. In each of these case, the entropy takes the same form as for the black hole horizon. Then, constructing a local causal horizon in every regular spacetime point, we can analyse the behaviour of their entropy. It turns out that the thermodynamic equilibrium conditions imposed on such local horizons encode (locally) the equations governing the gravitational dynamics. If the equivalence principle [^2] More precisely, one in general needs the Einstein equivalence principle. Specifically, to recover the Einstein equations a more stringent formulation, the strong equivalence principle (valid for gravitational test physics) is required [15]. For detailed discussion of the various formulations of the equivalence principle and their relations, see [33]. holds, these equations are valid throughout the spacetime. For entropy proportional to the horizon area, this approach recovers Einstein equations [4, 15, 19]. If the input of the derivation is any Wald entropy density [6] constructed from the metric and the Riemann tensor, it yields the equations of motion of the corresponding modified theory of gravity [16, 20, 21, 22]. Moreover, one can phrase the local equilibrium conditions entirely in terms of quantum von Neumann entropy. Then, they encode the semiclassical Einstein equations in which the quantum expectation value of the energy-momentum tensor serves as the source term for the classical spacetime curvature [17, 18, 19, 20].
@@
Herein, inspired by this success of thermodynamic methods in recovering the classical and semiclassical gravitational dynamics, we employ them to peer one step further. Namely, we study the thermodynamic perspective on the low energy quantum gravitational effects. We follow a research program we started some years ago, looking at the leading order quantum gravitational correction to entropy of causal horizon. It turns out to be proportional to the logarithm of the horizon area. Remarkably, a leading order correction term of this form is nearly universally predicted by various approaches to quantum gravity [34, 35, 36, 37, 38, 39] and also by some theory-independent considerations [40, 41, 42, 43, 44, 45, 46, 47, 48]. This generality of the logarithmic term in horizon entropy makes its influence on gravity robust and independent of any specific approach to quantum gravity. We have previously studied the effect of this term on the linearised gravitational dynamics. In this case, we found a result equivalent to the linearised equations of motion of quadratic gravity. We have also shown that beyond the linearised regime, the logarithmic contribution to entropy implies correction terms quadratic in the Ricci tensor. Notably, such terms do not appear in quadratic gravity, marking a clear departure from it. However, in both special cases we studied so far the corrections to gravitational dynamics only occur in the presence of matter. In the present work, we ask whether the Einstein equations are modified even in vacuum. We show that imposing the local Lorentz invariance prohibits any such modifications.
A more speculative option lies in breaking the local Lorentz invariance by introducing a preferred local direction of time. Then, the thermodynamic derivation yields a single scalar equation resembling the Hamiltonian constrain of general relativity. The source term of this equation is proportional to the Bel-Robinson tensor contracted with the local direction of time in all indices [49, 50]. Such a scalar expression is known as the Bel-Robinson super-energy density. The Bel-Robinson tensor is quadratic in the Weyl tensor and satisfies an analogy of the dominant energy condition valid for the energy-momentum tensor. It has been suggested that the Bel-Robinson super-energy density can be interpreted as a quasilocal gravitational energy density per unit surface area [51, 50]. This interpretation is consistent with the role the Bel-Robinson super-energy density plays in the geometry of small spheres, where it accounts for the area deficit due to spacetime curvature in vacuum [52, 53]. Then, its appearance as a source term for the Hamiltonian constraint is somewhat intuitive, as in general relativity the source is the matter-energy density. Moreover, it is well established that both the Bel-Robinson super-energy [52, 53, 54] and other measures of the gravitational energy [55, 56, 57] play a role in the dynamics of surfaces and, hence, in the gravitational thermodynamics. It is then to be expected that they can also act as source terms in the thermodynamically derived equations for gravitational dynamics. In total, while we leave the in depth exploration of this Lorentz invariance violating setup for future works, we find the possibility interesting enough to warrant a discussion in the present paper.
@@
Unless specified otherwise, we work in $D=4$ spacetime dimensions. We choose the mostly positive metric signature (,+,+,+)(-,+,+,+).To keep track of quantum and gravitational effects, we maintain $\hbar$ and $G$ explicit, but we set $c=k_{\text{B}}=1$. Other conventions follow [59].
@@ II.1 Light-cone cut local causal diamonds
We start by introducing the specific realisation of local, observer-dependent causal horizons we work with in this paper: the causal diamonds. In flat spacetime a causal diamond can be simply defined as the domain of dependence of a spacelike co-dimension $1$ ball. In a generic curved spacetime only small causal diamonds can be meaningfully defined. More specifically, one requires that the geodesic radius of the co-dimension $1$ ball forming the base of the causal diamond is much smaller than the local curvature length scale (i.e. the inverse of the square root of the largest eigenvalue of the Riemann tensor). Then, the spacetime curvature effects on the geometry of the causal diamonds can be treated as small corrections. However, even in this case, there exist several inequivalent generalisations of a causal diamond [58]. While we could in principle consider any of them, the definition best suited for our purposes is the light-cone cut local causal diamond, as we explain in the following.
The construction of a light-cone cut local causal diamond starts at an arbitrary regular spacetime point $A_{\text{p}}$, the past apex of the eventual causal diamond. At this point, we choose a unit timelike vector $n^{\mu}$ to represent a local direction of time. We then consider a class of future directed null vectors $k^{\mu}$ at $A_{\text{p}}$ normalised so that $n_{\mu}k^{\mu}=-1$, and define the past boundary of the light-cone cut local causal diamond as the congruence of null wordlines tangent to $k^{\mu}$. Any spacelike cross-section of this congruence is an approximate $2$-sphere (up to curvature-dependent corrections) and its interior an approximate $3$-dimensional spacelike ball. The base of the causal diamond corresponds to its spatial slice $\Sigma_{0}$ at the affine geodesic parameter length $l$. To construct the future (contracting) part of the light-cone cut local causal diamond, we introduce a future-directed null vector field $l^{\mu}$ defined on the boundary of $\Sigma_{0}$, so that $n_{\mu}l^{\mu}=-1$. Furthermore, denoting the projection of $k^{\mu}$ on the surface orthogonal to $n^{\mu}$ by $m^{\mu}$ (this vector is simply a unit, outward-pointing radial vector), we choose $l^{\mu}$ to have this projection equal to $m^{\mu}$. Then, the contracting congruence of the null geodesics tangent to $l^{\mu}$ makes up the future boundary of the causal diamond. We show a sketch of the light-cone cut local causal diamond in figure 1.
@@
The light-cone cut local causal diamonds then offer the additional advantage of fully specifying the null boundary of the causal diamond (up to some residual freedom [22, 25] we discuss in the following), while keeping the geometry of its spatial cross-sections somewhat ambiguous [58]. Then, it becomes very simple to track the physical matter entropy flux across the boundary and the corresponding changes in its area. Since, as we review in the following, the equilibrium conditions imposed on these two processes encode the equations governing the gravitational dynamics, light-cone cut local causal diamonds are ideally suited for deriving them in this way.
@@ II.2 Gravitational dynamics from thermodynamics
We have seen that the null boundary of a light-cone cut causal diamond represents a conformal Killing horizon. Then, it can be argued to possess entropy proportional to the area $\mathcal{A}$ of its spatial cross-section, i.e., $S=\eta\mathcal{A}$, where $\eta$ denotes the proportionality constant. This entropy can be interpreted as the von Neumann entropy sourced of the quantum entanglement between the two regions separated by the horizon [27, 28, 30]. However, more generally, it may also be understood as Shannon entropy accounting for the lack of information the observer inside the diamond has about its exterior [1]. Interested reader can find a more detailed discussion of this issue with further references here [25]. For the purposes of the present work, we do not care about the microscopic origin of this entropy, we simply take it as our assumption. The entropy, of course, changes as the area of the horizon expands. At the same time, matter flux crossing the horizon also carries entropy with it. It has been shown that the balance between both entropies encodes the gravitational dynamics [22, 25]. In the following, we review the main points of this argument, first recalling the expressions for the changes of both relevant entropies.
To compute the matter entropy flux, we employ the Clausius equilibrium relation $\Delta S_{\text{C}}=\Delta Q/T_{\text{U}}$. Here, $T_{\text{U}}$ denotes the Unruh temperature measured by accelerating observers moving inside the causal diamond and $\Delta Q$ the heat flux perceived by the same observer. The total flux of this Clausius entropy over the past boundary of the causal diamond equals (for details on its definition and computation, see [61])
@@
Since, by definition, the diamond expands to the past of the base $\Sigma_{0}$ and contracts to the future of it, the smooth evolution of the expansion requires $\theta_{(0)}=0$. It has been shown that non-zero $\sigma_{0,\mu\nu}$ leads to non-equilibrium entropy production [15, 63] analogous to the tidal heating of black holes [64]. This non-equilibrium entropy production can be straightforwardly dealt with, treating it in the same way as the $\theta_{\text{flat}}$ contribution we discuss in the following. Nevertheless, for simplicity, we set here $\sigma_{0\mu\nu}=0$. Without this technical condition, everything proceeds in the same way, but the higher order calculations in the next section become needlessly messy. Upon fixing $\sigma_{0\mu\nu}=0$, it is easy to see that the shear contributions to the expansion are not relevant and we obtain the following solution
@@
We can carry out the integration, using $k^{\mu}=n^{\mu}+m^{\mu}$ and the well known results for integrating unit radial vectors over a $2$-sphere
$\int m^{\mu}\text{d}\Omega_{2}$ $=0,$ (14)
$\int m^{\mu}m^{\nu}\text{d}\Omega_{2}$ $=\frac{4\pi}{3}\gamma^{\mu\nu}.$ (15)
where $\text{d}\Omega_{2}=\sin\theta\text{d}\theta\text{d}\phi$ is the angular integration element on the $2$-sphere and we introduced the spatial metric $\gamma^{\mu\nu}=g^{\mu\nu}+n^{\mu}n^{\nu}$. The result reads
@@
Equations (18) closely resemble the traceless Einstein equations. Taking their Newtonian limit allows us to identify the Newton gravitational constant $G$. Then, $G$ is defined in terms of the Planck constant $\hbar$ and the entropy proportionality constant $\eta$, i.e., $G=1/\left(4\hbar\eta\right)$. Then, we indeed recover the traceless Einstein equations [^3] Let us note that there exists non-metric, first order actions leading to dynamics equivalent to that given by equation (19), notably the Plebański-like action [65, 66] and the geometric trinity formulation of unimodular gravity [67]. Whether such first order dynamics can be recovered from thermodynamic equilibrium conditions remains an open question. Nevertheless, thermodynamics of first order theories is well established [68] and it should be possible to use it to recover the corresponding equations of motion.
@@
Let us conclude with two short remarks regarding the interpretation of equations (19). First, they are traceless, with the cosmological constant appearing only as an on-shell integration constant. Moreover, the geometry of the light-cone cut local causal diamonds is insensitive to the overall conformal factor of the metric. We have previously shown that these features imply that the gravitational dynamics we recover from thermodynamics (without any additional assumptions) does not correspond to general relativity, but rather to Weyl transverse (unimodular) gravity [23, 24, 25]. This theory has the same classical solutions as general relativity [69, 70, 71]. Moreover, the quantum dynamics of both theories have so far been found equivalent [71, 72, 73, 74]. The important difference is that in Weyl transverse gravity the cosmological constant is not sourced by the vacuum fluctuations, which makes its value radiatively stable [69, 75, 76, 77, 78, 79, 80]. Note that vacuum fluctuations not proportional to the metric tensors do gravitate, allowing for quantum gravitational corrections to the equations of motion of the kind described in the next section.
Therefore, the corrections to gravitational dynamics we discuss in the following should also be seen as modifying Weyl transverse gravity rather than general relativity.
Second, the fact that thermodynamics of local causal horizons recovers the equations governing gravitational dynamics is often interpreted as suggesting that gravity is emergent. In other words, the Einstein equations would be essentially equations of state describing the behaviour of some unknown quantum degrees of freedom of the spacetime in the classical limit [4, 14, 15, 81]. While we find this viewpoint worthwhile and thought provoking, we believe it is simply unnecessary to embrace it. We have shown that thermodynamics leads to novel insights into gravitational dynamics, even without requiring that the latter is emergent [23, 24, 25, 82]. All we need to assume is that the local equilibrium conditions encode all the information necessary to reconstruct the gravitational dynamics. We maintain this viewpoint in the rest of the paper.
@@ III Higher order corrections in vacuum: the Bel-Robinson tensor
In the previous section, we have reviewed the derivation of the traceless Einstein equations from thermodynamics. We neglected the subleading contributions to the curved spacetime changes of the area of the spatial cross-section of the horizon, and, hence, to its entropy. It is natural to ask whether these terms encode some corrections to Einstein equations. In some special cases, we have previously shown that this is indeed the case. However, such corrections only occur if we also take into account the subleading quantum contribution to the horizon entropy proportional to the logarithm of its area [82, 83]. Specifically, we have shown that the derivation generically leads to corrections to the traceless Einstein equations quadratic in the Ricci tensor [83, 84]. We have also derived the linearised modified dynamics, finding a result equivalent to a special case of linearised quadratic gravity [82]. Due to various technical and conceptual challenges, we are for the time unable to derive the modified equations in the fully general case. Here, we instead study another simplified setting, the vacuum case. On the one hand, we use this setting to develop tools that should allow us to derive the completely general equations in a subsequent work. On the other hand, any possible modifications to vacuum gravitational dynamics are of great interest by themselves, being the only kind of corrections relevant to the standard black hole solutions, in particular, to the Kerr metric (see [85, 86] for examples of such corrections). They are also important for the behaviour of the spacetime in the vicinity of generic spacelike singularities, as described by the Belinskii-Khalatnikov-Lifshitz conjecture [87]. Finally, such corrections generically modify the propagation of gravitational waves, which might make them experimentally falsifiable by the current or future gravitational waves detectors [88]. To study possible vacuum corrections to gravitational dynamics implied by the thermodynamic derivation, we can follow the same basic strategy we employed to derive the traceless Einstein equations. However, we need to introduce some conceptual changes, both on the level of the initial assumptions and in the process of the derivation. Therefore, we start by reviewing the requirements we impose.
@@
In the semiclassical setup, we have relied on the Einstein equivalence principle to construct the local Minkowski vacuum necessary for invoking the Unruh effect. However, the status of any formulation of the equivalence principle in a regime in which quantum gravitational corrections become relevant remains an open issue. Therefore, we now instead impose two more modest requirements, which do not rely on the general formulation of the equivalence principle but still allow for a thermodynamic description of local causal horizons. Our first requirement then concerns an observer equipped with a particle detector and moving at an approximately constant acceleration. Given enough time for the detector to thermalise, we require that it registers local Minkowski vacuum as a thermal bath of particles at the Unruh temperature $T_{\text{U}}=a/\left(2\pi\right)$. While the fate of the Unruh effect in quantum gravity remains an open questions, there exist indications that it remains viable [89].
@@
The second requirement we impose is a modified entropy which includes the leading order quantum corrections including corrections from quantum contributions [^4] Note here, that it is possible to explore more radical modifications of entropy, including non-extensive proposals, and study the possible modified gravitational dynamics they encode. While this option constitutes a natural extension of our research programme, we do not pursue it here.. In particular, entropy of a spherically symmetric local causal horizon in four spacetime dimensions has a generic form
@@
where $\mathcal{A}_{0}$ is a constant with the dimensions of area and $\mathcal{C}$ a real number. Both $\mathcal{A}_{0}$ and $\mathcal{C}$ are characteristic for the particular approach to computing entropy, but independent of the horizon for which it is computed. The first two terms in equation (20) remain in the limit $l_{\text{P}}\to 0$ (and the leading one diverges), whereas the $O\left(\mathcal{A}_{0}/\mathcal{A}\right)$ contributions vanish. Then, restricting our attention to the area term and the logarithm is equivalent to keeping the only two contributions to entropy relevant for large horizons, such that $\mathcal{A}\gg l_{\text{P}}^{2}$.
Ideally, parameter $\mathcal{C}$ could be measured in some experimental set up to detect quantum gravity corrections. However, like other quantum gravitational corrections, the value of C is beyond the current experimental limits. Nevertheless, there already exist proposals to establish bounds on $\mathcal{C}$ [92] and quantify the changes this parameter implies, e.g. for the thermal spectrum of particles produced by the Hawking effect [93].
Equation (20) arises independently in a number of different contexts. It appears, e.g. in loop quantum gravity [34, 35], string theory [36, 37, 38], AdS/CFT correspondence [39], entanglement entropy calculations [30, 41], phenomenological approaches based on the existence of a minimal length [42, 43], in an analysis of statistical fluctuations [44, 45, 46] and from quantisation of the horizon area [47, 48]. In other words, the logarithmic correction to entropy associated with causal horizons is present in most of the approaches to quantum gravity. The specific approach is characterised just by the sign and value of the parameter $\mathcal{C}$. Thence, the modifications of gravitational dynamics implied by the logarithmic term are in principle relevant for any approach to quantum gravity. They are only parametrised by a single unknown number, $\mathcal{C}$, which is of the order of unity for most approaches.
@@
We now need to solve the Raychaudhuri equation (5) and equation (10) for the evolution of the shear tensor to the order $O\left(\lambda^{3}\right)$ (recall that the affine null geodesic parameter $\lambda$ lies in the interval from l-l to 0 on the past boundary), which becomes relevant once we take into account the logarithmic corrections to entropy. We obtain (keeping in place the condition $\theta_{0}=\sigma_{0\mu\nu}=0$)
$\theta\left(x\right)=$ $\theta_{\text{flat}}\bigg[1+\frac{\lambda^{2}}{1+\lambda\theta_{\text{flat}}}R_{\mu\nu}k^{\mu}k^{\nu}+\frac{1}{6}\frac{2+\lambda\theta_{\text{flat}}}{\left(1+\lambda\theta_{\text{flat}}\right)^{2}}\lambda^{4}\left(R_{\mu\nu}k^{\mu}k^{\nu}\right)^{2}\bigg]$